OpenStudy (anonymous):

Solve the equation for t:

OpenStudy (anonymous):

\[7000e ^{.09t}=20000(1+.033)^{t}\]

OpenStudy (anonymous):

first divide by 1000 and take the ln of both sides (natural logarithm) ln7e^.09t=ln20(1.033)^t

OpenStudy (anonymous):

you have to remember the log rules ln7+0.9t=ln20+t*ln1.033

OpenStudy (anonymous):

from this you just express t t(0.9-ln1.033)=ln20-ln7 t=(ln20-ln7)/(0.9-ln1.033)

OpenStudy (anonymous):

that is actually wrong..

OpenStudy (anonymous):

Nevermind the .9 should just be .09

OpenStudy (anonymous):

Nevermind again that answer is still wrong!